### v1

# Questioning Cosmic Acceleration with DESI: The Big Stall of the Universe

Abstract
One of the most important discoveries in modern cosmology is cosmic acceleration. However, we find that today’s universe could decelerate in the statistically preferred Chevallier-Polarski-Linder (CPL) scenario over the $\Lambda$CDM model by cosmic microwave background, type Ia supernova and DESI’s new measurements of baryon acoustic oscillations. Using various datasets, at a beyond $5\,\sigma$ confidence level, we demonstrate that the universe experiences a triple deceleration during its evolution and finally reaches the state of the “Big Stall”, which predicts that: (i) the universe suddenly comes to a halt in the distant future; (ii) its eventual destiny is dominated by dark matter rather than dark energy ; (iii) it ultimately retains an extremely small fraction of dark energy but exerts an extremely large pressure. Our findings profoundly challenge the established understanding of cosmic acceleration and enrich our comprehension of cosmic evolution.
Introduction. The discovery of cosmic acceleration stands as one of the most profound revelations in modern cosmology, fundamentally altering our understanding of the universe’s evolution and destiny. This phenomenon, which depicts the observed increase in the rate of the universe’s expansion, has captivated scientists since its unexpected revelation. The implications of cosmic acceleration are vast, touching upon the very nature of space, time, and the fundamental forces that govern the cosmos.
Historically, the prevailing cosmological model suggested that the universe’s expansion, initiated by the Big Bang, would gradually decelerate over time due to the gravitational attraction of matter. This view was grounded in Einstein’s theory of general relativity and supported by observations of the early cosmic microwave background (CMB) radiation and the large-scale structure of the universe COBE:1992syq ; Peebles1993 ; Mather:1993ij ; Bennett:1996ce . However, precise measurements of distant Type Ia supernovae (SN) challenged this perspective in the late 1990s. Two independent teams, the High-z Supernova Search Team SupernovaSearchTeam:1998fmf and the Supernova Cosmology Project SupernovaCosmologyProject:1998vns , revealed that these stellar explosions appeared fainter than expected. This inconsistency implied that cosmic expansion was not slowing down but rather accelerating.
The concept of cosmic acceleration introduced a new component into the cosmological framework: dark energy (DE). This enigmatic force, which permeates the fabric of space itself, is believed to drive the accelerated expansion of the universe. Nowadays, the standard cosmological model is the so-called $\Lambda$CDM, where DE is usually represented by the cosmological constant ($\Lambda$). It also includes cold dark matter (CDM) and ordinary baryonic matter Bahcall:1999xn ; Turner:1997npq ; Carroll:2000fy . While the $\Lambda$CDM scenario has been remarkably successful in explaining a wide range of cosmological observations, the nature of DE remains one of the most profound mysteries in contemporary physics Weinberg:1988cp ; Peebles:2002gy ; Padmanabhan:2002ji ; Frieman:2008sn .
Understanding cosmic acceleration is not merely an academic pursuit; it is essential for addressing fundamental questions about the universe’s ultimate fate. If the acceleration continues unabated, the universe will expand indefinitely, leading to a cold and desolate future known as the “heat death” of the universe Dyson:1979zz ; Adams:1996xe . In this scenario, the universe becomes increasingly empty as galaxies move farther apart, stars burn out, and the cosmos approaches a state of maximum entropy. This fate implies a universe devoid of usable energy, where no further work can be done, and all physical processes eventually cease.
Alternatively, if the nature of DE evolves over time, the universe’s fate could be dramatically different. One possibility is the “Big Rip”, where the accelerating expansion becomes so rapid that it overcomes the gravitational, electromagnetic, and nuclear forces holding matter together Caldwell:2003vq . In this scenario, galaxies, stars, and even atoms are torn apart, leading to a catastrophic end of the universe. Another intriguing possibility is a cyclical model of the universe, where periods of expansion and contraction repeat indefinitely Steinhardt:2002ih ; Penrose:2010zz . In such models, the universe undergoes cycles of birth, expansion, contraction, and rebirth, potentially driven by the dynamics of DE or modifications to general relativity.
Furthermore, some theories propose that the accelerated expansion might slow down or even reverse in the distant future, leading to a “Big Crunch” where the universe collapses back into a singularity Barrow:1986 . Other speculative scenarios include the “Big Freeze”, similar to the heat death but emphasizing the extreme cold and emptiness of the future universe, and the “Vacuum Decay”, where a phase transition in the vacuum energy leads to a sudden and dramatic change in the properties of the universe Coleman:1980aw .
During the past two decades, the CMB Planck:2018vyg ; ACT:2025fju ; SPT-3G:2022hvq ; WMAP:2003elm ; Planck:2013pxb , baryon acoustic oscillations (BAO) SDSS:2005xqv ; 2dFGRS:2005yhx ; Beutler:2011hx ; eBOSS:2020yzd ; deCarvalho:2017xye ; eBOSS:2017cqx , weak gravitational lensing Heymans:2012gg ; Hildebrandt:2016iqg ; Planck:2018lbu ; DES:2017qwj , galaxy clustering DES:2017myr ; DES:2021wwk , cluster counts DES:2025xii and SN observations SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns have further deepened our understanding of cosmic acceleration and confirmed the validity of $\Lambda$CDM. However, it confronts at least two challenges, i.e., the cosmological constant conundrum Weinberg:1988cp ; Carroll:2000fy ; Peebles:2002gy ; Padmanabhan:2002ji and the coincidence problem Steinhardt:1997 , while suffering from the emergent cosmic tensions of the Hubble constant ($H_{0}$) and the matter fluctuation amplitude ($S_{8}$) DiValentino:2020vhf ; DiValentino:2020zio ; DiValentino:2020vvd ; Abdalla:2022yfr ; DiValentino:2025sru . Logically, these advancements compel theorists to study new physics beyond the $\Lambda$CDM model in order to resolve these issues (see Abdalla:2022yfr ; DiValentino:2025sru for reviews).
Recently, the new BAO measurements from the DESI data release two (DR2) DESI:2024mwx ; DESI:2024uvr ; DESI:2024lzq ; DESI:2025zgx ; DESI:2025zpo inspired the explorations of new physics once again, especially, dynamical dark energy (DDE) DESI:2025fii . Through the combinations of CMB, BAO and SN data, the DESI collaboration reported substantial evidences of DDE. However, in Wang:2025bkk , we question the validity of such combinations, because there exist clear tensions among the datasets. Currently, a reliable approach is constraining the evolution of DE with each independent dataset and observe if there are statistical evidences of DDE. Interestingly, we find strong preferences of DDE over $\Lambda$CDM for each dataset used. This raises an even more interesting question: Whether cosmic acceleration still holds in such a preferred DDE universe? Employing various datasets, we find that current individual datasets cannot determine whether the present-day universe is accelerating or not.
Figure 1: One-dimensional and two-dimensional posterior distributions of the parameter pairs $\Omega_{m}$-$\omega_{0}$ (left) and $\omega_{0}$-$\omega_{a}$ (right) from various datasets in the CPL model. The chocolate and olive solid lines are the boundaries of cosmic acceleration and super acceleration. The green dashed lines represents $\Omega_{m}=0.5$ (left) and $\omega_{a}=0$ (right). The red and blue dashed lines denote $\omega_{0}=0$ and $\omega_{0}=-1$, respectively.
Table 1: Mean values and $1\,\sigma$ (68.3%) uncertainties of the present-day deceleration parameter $q_{0}$ from different datasets in the CPL model.
       Data        DESI        CMB        Pantheon+        Union3        DESY5
       $q_{0}$        $0.34\pm 0.41$        $2.83\pm 1.89$        $-0.35^{+0.20}_{-0.21}$        $0.12^{+0.29}_{-0.27}$        $0.19\pm 0.28$
Figure 2: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from five individual datasets in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.
Model. In the framework of general relativity Einstein:1916vd , considering a homogeneous and isotropic universe, the Friedmann equations read as $H^{2}=(8\pi G\rho)/3$ and $\ddot{a}/a=-4\pi G(\rho+3p)/3$, where $H$ is the cosmic expansion rate at a scale factor $a$ and $\rho$ and $p$ are the mean energy density and pressure of different species including baryons, dark matter (DM) and DE in the late universe. Combining two Friedmann equations, the normalized Hubble parameter $E(z)\equiv H(z)/H_{0}$ for a flat CPL universe Chevallier:2000qy ; Linder:2002et is shown as
$E(a)=\left[\Omega_{m}a^{-3}+(1-\Omega_{m})a^{-3(1+\omega_{0}+\omega_{a})}\mathrm{e}^{3\omega_{a}(a-1)}\right]^{\frac{1}{2}},$ (1)
where $\Omega_{m}$ is the matter fraction. It reduces to $\Lambda$CDM when $\omega_{0}=-1$ and $\omega_{a}=0$. In terms of the deceleration parameter $q\equiv-\ddot{a}/(aH^{2})$, cosmic acceleration requires $\omega_{0}<1/[3(\Omega_{m}-1)]$ easily derived from the violation of strong energy condition $\rho+3p\geqslant 0$, while a super acceleration requires $\omega_{0}<1/(\Omega_{m}-1)$ corresponding to the violation of null energy condition $\rho+p\geqslant 0$.
[…]
We use the Boltzmann code CAMB Lewis:1999bs to compute the background evolution and theoretical power spectra of the universe. To perform the Bayesian analysis, we take the Monte Carlo Markov Chain (MCMC) method to infer the posterior distributions of model parameters using Cobaya Torrado:2020dgo . We assess the convergence of MCMC chains via the Gelman-Rubin criterion $R-1\lesssim 0.01$ Gelman:1992zz and analyze them using Getdist Lewis:2019xzd .
We take the uniform priors for free parameters: the baryon fraction $\Omega_{b}h^{2}\in[0.005,0.1]$, cold DM fraction $\Omega_{c}h^{2}\in[0.001,0.99]$, acoustic angular scale at the recombination epoch $100\theta_{\rm MC}\in[0.5,10]$, scalar spectral index $n_{s}\in[0.8,1.2]$, amplitude of the primordial scalar power spectrum $\ln(10^{10}A_{s})\in[2,4]$, optical depth $\tau\in[0.01,0.8]$, present-day DE EoS $\omega_{0}\in[-15,20]$ and amplitude of DE evolution $\omega_{a}\in[-30,10]$. To produce a matter-dominated era at high redshifts, we impose the condition $\omega_{0}+\omega_{a}<0$ in the Bayesian analysis. The reason why we use such wide priors for ($\omega_{0}$, $\omega_{a}$) is that a large enough parameter space can completely exhibit the constraining power of DESI BAO data.
Deceleration preferred by data. Although the Planck collaboration has given a $2\,\sigma$ evidence of DDE Planck:2018vyg , enough attention is not paid to it due to very loose constraints on ($\omega_{0}$, $\omega_{a}$) and the reason that combining previous BAO and SN with CMB data gives no deviation from $\Lambda$CDM Planck:2018vyg . In light of the fact that new DESI BAO and SN data independently prefer DDE over $\Lambda$CDM Wang:2025bkk , the CMB DDE evidence should be given sufficient attention. In Fig.1, we present the parameter spaces of ($\Omega_{m}$, $\omega_{0}$) from eight independent datasets including CC, GRB and HII observations, and find that, except for Pantheon+ that give a $1.75\,\sigma$ hint of cosmic acceleration, most of the parameter space from each dataset supports cosmic deceleration (see mean values in Tab.1). Specifically, CMB provides a $1.5\,\sigma$ evidence for deceleration, while DESI, Union3 and DESY5 gives $\sim 1\,\sigma$, $0.43\,\sigma$ and $0.68\,\sigma$ clues, respectively. To demonstrate the stability of our findings, using eight datasets, we show constraints on ($\omega_{0}$, $\omega_{a}$) in Fig.1. It is easy to see that all the datasets prefer the DDE region of $\omega_{0}>-1$ and $\omega_{a}<0$. Note that the CPL model is statistically favored over $\Lambda$CDM Wang:2025bkk for these datasets. Therefore, we have verified that the universe could possibly undergo a present-day deceleration in such a preferred DDE universe by data. Interestingly, all the datasets cannot rule out $\omega_{0}>0$, indicating the possibility that the pressure of DE is positive.
Furthermore, employing the error propagation, we derive the redshift evolution of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from different datasets in the CPL model (see Fig.2). Overall, all the datasets support a triple deceleration at the $5\,\sigma$ confidence level, although the redshift ranges of deceleration are slightly different. It is a phenomenon similar to the staged acceleration of a rocket, but in the form of staged deceleration. One phase of deceleration occurred in the past, while the remaining two phases are anticipated to occur in the future. Unfortunately, we cannot determine if today’s universe is accelerating with a high significance.
The Big Stall. The $\Lambda$CDM model suggests a future where the universe becomes increasingly cold, dark, and empty, with the accelerating expansion driven by DE leading to the ultimate isolation and decay of all structures. However, this is not the case in the CPL model. In Fig.2, we find that all five datasets do not exclude the possibility of present-day matter domination, although DESI DR2 gives a $2.35\,\sigma$ evidence of $\Omega_{m}<0.5$. Based on five individual datasets, we confirm the following cosmic expansion history: After a long-term deceleration sourcing from the matter domination, DE starts to become comparable in magnitude to matter and mitigates the rate of deceleration of the universe. However, we cannot ensure whether the universe is accelerating now. Then, the universe will experience the second deceleration when the DE fraction significantly decreases and the universe becomes matter-dominated in the future. As the deceleration of the universe continues decreasing, it will enter the second phase where it is uncertain whether acceleration will resume. Subsequently, when the matter fraction is large enough, the universe will experience continuous deceleration until its acceleration ultimately reaches zero at the end of its existence. Interestingly, throughout this process, the rate of cosmic expansion continuously decreases and also reaches zero at the end of the universe. Conversely, the DE EoS continue increasing and finally become a positive infinity. This indicates that the ultimate fate of the universe is completely stationary with $q(z\rightarrow-1)\rightarrow 0$ and $E(z\rightarrow-1)\rightarrow 0$ and matter-dominated. In the meanwhile, the universe will end up with an extremely small DE fraction but an extremely large DE pressure. We call this fate as the “Big Stall”. Using the information of the baryon fraction from CMB Planck:2018vyg or BBN Cooke:2017cwo , the universe is DM-dominated at its end. Note that all the datasets roughly reveal that DE exhibits a phantom-crossing behavior at $2\,\sigma$ level, i.e., $\omega_{0}<-1$ in the past and $\omega_{0}>-1$ (and $\omega_{0}>0$) in the future.
Concluding remarks. Currently, we cannot determine whether the universe is accelerating and even CMB, DESI DR2, Union3 and DESY5 data slightly prefer cosmic deceleration. This is different from the claim made by two SN teams who confirm cosmic acceleration under $\Lambda$. Because all the individual datasets favor the CPL DDE region of $\omega_{0}>-1$ and $\omega_{a}<0$ and prefer the CPL scenario over $\Lambda$CDM by a statistical comparison Wang:2025bkk , our findings are logically reasonable. So far, there is no any independent dataset that can demonstrate the existence of today’s cosmic acceleration in such a preferred universe. Based on this concern, the important contribution from two SN teams should be discovering the existence of dark energy not cosmic acceleration SupernovaSearchTeam:1998fmf ; SupernovaCosmologyProject:1998vns .
The combinations of CMB, DESI and SN data give a similar fate of the universe to individual datasets. The main difference is that they strongly support today’s cosmic acceleration or DE domination. However, their results are problematic DESI:2024mwx ; DESI:2025zgx , since they suffer from tensions among the datasets. For completeness, we show their background quantities in the supplementary material.
In the CPL model, current observations give a beyond $5\,\sigma$ evidence of the Big Stall with a triple deceleration for the fate of the universe. Completely different from the Big Freeze and the Big Rip, the Big Stall predicts the following novel properties about the fate of the universe: (i) continuously decelerate; (ii) suddenly halt at a distant future; (iii) DM-dominated; (iv) extremely small DE fraction with extremely large pressure and EoS; (v) stars and galaxies could still be active; (vi) galaxies are not isolated; (vii) no black hole era from the Big Freeze; (viii) it will not reaches a state of maximum entropy.
Acknowledgements. DW is supported by the CDEIGENT fellowship of Consejo Superior de Investigaciones Científicas (CSIC).
Appendix A Supplementary Material for “Questioning Cosmic Acceleration with DESI: The Big Stall of the Universe”
     In this supplementary material, we exhibit the redshift evolution of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from four data combinations including CMB+DESI, CMB+DESI+Pantheon+, CMB+DESI+Union3 and CMB+DESI+DESY5 in the CPL model (see Fig.1).
     Overall, these combinations provide tighter constraints on each background quantity due to the degeneracy breaking. All the combinations support today’s universe is accelerating and dominated by dark energy, while they give stronger evidences of $\omega_{0}>-1$ (and $\omega_{0}>0$) in the distant future. Very interestingly, same as individual datasets, all these combinations reveal, a beyond $5\,\sigma$ confidence level, that the universe experiences a triple deceleration during its evolution and finally reaches the state of the “Big Stall”. Therefore, the beyond $5\,\sigma$ signal of the universe with a triple deceleration and the Big Stall does not depend on which datasets are used.
Figure 1: The $1\,\sigma-5\,\sigma$ regions of background quantities $\Omega_{m}(z)$, $q(z)$, $E(z)$ and $\omega(z)$ from different data combinations in the CPL model. As a comparison, we plot $q(z)$ on a logarithmic scale in the third row. The dashed lines denote the mean values of each quantity (red), $z=0$ (blue), $\Omega_{m}=0.5$ (black), $q=0$ (green), $E=1$ (purple), $\omega_{0}=0$ (orange) and $\omega_{0}=-1$ (pink), respectively.